# Schur–Weyl Equivalences for Wreath Product Superalgebras

### v1

@@ 3.3. Defining a $S_{d}\wr A$-module structure
be the function which counts the number of differences in the entries of $(\mathbf{i},\mathbf{i})_{k}$ versus the entries of $((1,2,\dotsc,d),(1,2,\dotsc,d))_{k}$. We argue by induction on $D\left((\mathbf{i},\mathbf{i})_{k}\right)$, with the base case of $D\left((\mathbf{i},\mathbf{i})_{k}\right)=0$ being trivial.
Assume $D\left((\mathbf{i},\mathbf{i})_{k}\right)>0$. Then $i_{\ell}\neq\ell$ for some $\ell\in[1,d]$. We first consider the case when $\ell$ does not appear as an entry of $(\mathbf{i},\mathbf{i})_{k}$. If $\ell\neq k$, then we may use (3) from the previous lemma to simultaneously change $i_{\ell}$ and $j_{\ell}$ to $\ell$ and thereby obtain a pair $(\mathbf{i}^{\prime},\mathbf{i}^{\prime})_{k}\in\tilde{I}^{2}(n,d)$ which is equivalent to $(\mathbf{i},\mathbf{i})_{k}$ and with $D((\mathbf{i}^{\prime},\mathbf{i}^{\prime})_{k})=D((\mathbf{i},\mathbf{i})_{k})-1$. If $\ell=k$, let $\mathbf{j}^{\prime}=\mathbf{i}^{\prime}=(i_{1},\dotsc,i_{k-1},\ell,i_{k+1},\dotsc,i_{d})$. Using (1) shows that $\alpha^{a}_{(\mathbf{i},\mathbf{i})_{k}}=\alpha^{a}_{(\mathbf{j}^{\prime},\mathbf{i})_{k}}$ and (2) shows $\alpha^{a}_{(\mathbf{j}^{\prime},\mathbf{i})_{k}}=\alpha^{a}_{(\mathbf{j}^{\prime},\mathbf{i}^{\prime})_{k}}$. Once again, we have obtained a pair $(\mathbf{i}^{\prime},\mathbf{i}^{\prime})_{k}\in\tilde{I}^{2}(n,d)$ which is equivalent to $(\mathbf{i},\mathbf{i})_{k}$ and with $D((\mathbf{i}^{\prime},\mathbf{i}^{\prime})_{k})=D((\mathbf{i},\mathbf{i})_{k})-1$.
@@
Thus, by induction, $(\mathbf{i},\mathbf{i})_{k}$ is equivalent to $((1,2,\dotsc,d),(1,2,\dotsc,d))_{k}$. This proves the claim. ∎
In light of the previous result, we sometimes write $\alpha^{a}_{k}$ for $\alpha^{a}_{(\mathbf{j},\mathbf{i})_{k}}$, where the $(\mathbf{j},\mathbf{i})_{k}\in\tilde{I}^{2}(n,d)$ is left implicit. We continue to assume that $N$ is a $\mathds{k}S_{d}$-supermodule such that $E_{n}^{\otimes d}\otimes_{\mathds{k}S_{d}}N$ has a $U(\mathfrak{gl}_{n}(A))$-supermodule structure that makes it a supermodule in $\mathcal{P}_{d}$, and such that the action of the subalgebra $U(\mathfrak{gl}_{n}(\mathds{k}))$ on $E_{n}^{\otimes d}\otimes_{\mathds{k}S_{d}}N$ is the obvious one.
@@
We first observe that $\alpha_{k}^{1}=\operatorname{Id}_{N}$ for all $k\in[1,d]$. Namely, fix $(\mathbf{j},\mathbf{i})_{k}\in\tilde{I}^{2}(n,d)$ and then
@@
From Theorem 3.1.1, $\alpha_{k}^{1}(n)=n$, as claimed.
@@
for all $k\in[1,d]$ and all homogeneous $a\in A$. Since the various $a_{k}:=1^{\otimes(k-1)}\otimes a\otimes 1^{\otimes(d-k)}$ generate $A^{\otimes d}$, it suffices to verify:
@@
We next explain the fourth claim. Pick $(\mathbf{j},\mathbf{i})_{k},(\mathbf{i},\mathbf{i})_{k}\in\tilde{I}^{2}(n,d)$ with $i_{k}$ and $j_{k}$ distinct. Two such pairs exist because $n\geq d+1$. We wish to show $\alpha_{(\mathbf{j},\mathbf{i})_{k}}^{a}\circ\alpha_{(\mathbf{i},\mathbf{i})_{k}}^{a^{\prime}}=\alpha_{(\mathbf{j},\mathbf{i})_{k}}^{aa^{\prime}}$. Using the commutator formula, we have
@@
However, by the second property of supermodules in $\mathcal{P}_{d}$, the first term in the last sum equals zero. Theorem 3.1.1 shows that $\alpha_{(\mathbf{j},\mathbf{i})_{k}}^{a}(\alpha_{(\mathbf{i},\mathbf{i})_{k}}^{a^{\prime}}(n))=\alpha_{(\mathbf{j},\mathbf{i})_{k}}^{aa^{\prime}}(n)$, as desired.
@@
Fix $(\mathbf{j},\mathbf{i})_{p}\in\tilde{I}^{2}(n,d)$. Set $\mathbf{j}^{\prime}=(j_{1},\dotsc j_{p-1},j_{p+1},j_{p},j_{p+2},\dotsc,j_{d})$ and $\mathbf{i}^{\prime}=(i_{1},\dotsc i_{p-1},i_{p+1},i_{p},i_{p+2},\dotsc,i_{d})$. Then $(\mathbf{j}^{\prime},\mathbf{i}^{\prime})_{p+1}\in\tilde{I}^{2}(n,d)$. Under the right place permutation action of $S_{d}$ on $E_{n}^{\otimes d}$, we have $\mathbf{v}_{\mathbf{j}^{\prime}}\sigma_{p}=\mathbf{v}_{\mathbf{j}}$ and $\mathbf{v}_{\mathbf{i}}\sigma_{p}=\mathbf{v}_{\mathbf{i}^{\prime}}$. We then calculate:
@@
Theorem 3.1.1 implies $\sigma_{p}(\alpha_{p}^{a}(\sigma_{p}(n)))=\alpha^{a}_{p}(n)$, as desired. ∎
@@ 3.4. Recovering the action of $\mathfrak{gl}_{n}(A)$
To summarize, if $n\geq d+1$ and $M$ is a $U(\mathfrak{gl}_{n}(A))$-supermodule in $\mathcal{P}_{d}$, then we have established the existence of a $S_{d}\wr A$-supermodule $N$ so that $E_{n}^{\otimes d}\otimes_{\mathds{k}S_{d}}N$ is isomorphic to $M$ as $U(\mathfrak{gl}_{n}(\mathds{k}))$-supermodules. Via this isomorphism there is an action of $U(\mathfrak{gl}_{n}(A))$ on the superspace $E_{n}^{\otimes d}\otimes_{\mathds{k}S_{d}}N$. As we next explain, one can explicitly describe the action of $\mathfrak{gl}_{n}(A)$ in terms of the maps $\alpha^{a}_{k}:N\to N$.
@@
Let $n\geq d+1$. Let $E_{n}^{\otimes d}\otimes_{\mathds{k}S_{d}}N$ be as above and let $n\in N$. For homogeneous $a\in A$, $k\in[1,d]$, $p,q\in[1,n]$, and $\mathbf{i}\in I(n,d)$, we have
@@
The base case is when all entries of $\mathbf{i}$ are distinct and $D(\mathbf{i})=0$. If $q$ does not appear as an entry of $\mathbf{i}$, then the left side of (3.4.1) is zero because $\boldsymbol{\varepsilon}_{\mathbf{i}}+\varepsilon_{p}-\varepsilon_{q}$ is not an element of $\Lambda^{+}(n,d)$, and the right side is also zero because $\delta_{q,i_{k}}=0$ for all $k$. If $q$ appears as an entry of $\mathbf{i}$, then it appears only once — say at position $\ell\in[1,d]$. Let $\mathbf{j}=(i_{1},\dotsc,i_{\ell-1},p,i_{\ell+1},\dotsc,d)$. In this case, the left hand side equals $\mathbf{v}_{\mathbf{j}}\otimes\alpha^{a}_{(\mathbf{j},\mathbf{i})_{\ell}}(n)$ and, by direct calculation the right hand side also equals $\mathbf{v}_{\mathbf{j}}\otimes\alpha^{a}_{\ell}(n)$. In both cases (3.4.1) is verified when $D(\mathbf{i})=0$ for all $E^{a}_{p,q}$.
We now assume $D(\mathbf{i})>0$ and that (3.4.1) holds for for all $E^{a}_{p,q}$ and all $v_{\mathbf{j}}\otimes n$ whenever $D(\mathbf{j})<D(\mathbf{i})$. Since $D(\mathbf{i})>0$, there exists $1\leq k_{1}<k_{2}\leq d$ such that $i_{k_{1}}=i_{k_{2}}$. Let $\ell=k_{1}$. Since $n\geq d+1$, we may choose $x\in[1,n]$ so that $x$ does not appear as an entry of $\mathbf{i}$. Set $\mathbf{j}=(i_{1},\dotsc,i_{\ell-1},x,i_{\ell+1},\dotsc,i_{d})$. Observe that $D(\mathbf{j})<D(\mathbf{i})$. From the supercommutator formula, we have
@@ 3.6. Main Theorem
When $n\geq d+1$, $F^{A}_{n,d}(N)\cong E_{n}^{\otimes d}\otimes_{\mathds{k}S_{d}}N$ is finite-dimensional if and only if $N$ is finite-dimensional. Thus, the restriction of $F^{A}_{n,d}$ defines an equivalence between the full subcategories of finite-dimensional supermodules.
@@ 3.7. Some Consequences
We give a few example applications of the main theorem. Throughout this section, we assume $A$ is finite-dimensional and that $\mathds{k}$ is a splitting field for $A$.
@@
Write $J(A)$ for the Jacobson radical of the superalgebra $A$. By [4, Lemma 2.6] this is a homogeneous ideal and coincides with the Jacobson radical of AA as an algebra. The arguments of [10, Section 3] show that
@@
By [10, Section 6], if $A$ is quasi-hereditary algebra, then so is $S_{d}\wr A$ and, hence, the category of $S_{d}\wr A$-modules is a highest weight category. Thus, we have the following result.
@@
Say $A$ finite-dimensional $\mathds{k}$-superalgebra and that $S=\left\{S_{1},\dotsc,S_{t}\right\}$ is a complete, irredundant set of simple $A$-supermodules up to isomorphism. Recall that a simple $A$-supermodule may either be of type M or of type Q. It is known that the simple $S_{d}\wr A$-supermodules are indexed by the set
@@
For example, see [22] for the construction of the simples when $A$ is an algebra or [25, 11] when $A$ is a superalgebra. Write $S^{\gamma}$ for the simple $S_{d}\wr A$-supermodule labelled by $\gamma\in\Gamma^{d}(A)$. We note that the type of $S^{\gamma}$ can be determined from the function $\gamma$ and that $F^{A}_{n,d}$ preserves the type of simple supermodules. Since we do not use this information, we omit the details.
@@
$\left\{L_{n}^{A}(\gamma):=F^{A}_{n,d}(S^{\gamma})\mid\gamma\in\Gamma^{d}(A)\right\}$
@@ 3.8. Sharpness of the Bound
Given that the equivalence of categories holds when $n\geq d$ in the classical (i.e., $A=\mathds{k}$) case, we could ask if the bound $n\geq d+1$ here is an artifact of the proof. The following example shows this is not the case. A similar claim was made for $A=\mathds{k}[t^{\pm 1}]$ in [13].
@@
Let $A=\mathds{k}[x]/(x^{2}-1)$ be the group algebra of the cyclic group of order two. Let $n=d$. We can inflate the sign representation of $\mathds{k}S_{d}$ through the canonical surjective map $S_{d}\wr A\to\mathds{k}S_{d}$ to define a one-dimensional simple $S_{d}\wr A$-module, $S=\mathds{k}u$.
Let
$L=F^{A}_{n,d}(S)=V_{n}^{\otimes d}\otimes_{S_{d}\wr A}S.$
Then $L$ is one-dimensional and spanned by $\mathbf{v}_{\mathbf{i}}^{\mathbf{1}}\otimes u$ where $\mathbf{i}=(1,\dotsc,n)$. To be explicit, using the notation established in Example 2.4.2, the action of $\mathfrak{gl}_{n}(A)$ is given by: $x(\mathbf{v}_{\mathbf{i}}^{\mathbf{1}}\otimes u)=0$ for $x\in\mathfrak{sl}_{n}(A)$, $I_{n}(\mathbf{v}_{\mathbf{i}}^{\mathbf{1}}\otimes u)=n(\mathbf{v}_{\mathbf{i}}^{\mathbf{1}}\otimes u)$, and $xI_{n}(\mathbf{v}_{\mathbf{i}}^{\mathbf{1}}\otimes u)=\mathbf{v}_{\mathbf{i}}^{\mathbf{1}}\otimes u$.
Observe that $L$ is a polynomial representation of $U(\mathfrak{gl}_{n}(\mathds{k}))$ of degree $d=n$. The spanning vector has weight $\boldsymbol{\varepsilon}_{\mathbf{i}}=\varepsilon_{1}+\dotsb+\varepsilon_{n}$ and, since $\boldsymbol{\varepsilon}_{\mathbf{i}}(h_{i})=1$ for $i=1,\dotsc,n$, this module vacuously satisfies the second condition of Definition 2.4.1. Thus, $L$ is in $\mathcal{P}_{d}(\mathfrak{gl}_{n}(A))$.
@@
Observe that $T$ restricted to $U(\mathfrak{gl}_{n}(\mathds{k}))$ is a polynomial representation of degree $n=d$. It again vacuously satisfies the second condition of Definition 2.4.1. Therefore, $T$ is a non-split extension of $L$ with itself in $\mathcal{P}_{d}(\mathfrak{gl}_{d}(A))$. On the other hand, $S_{d}\wr A$ is a semisimple algebra. Therefore, $\mathcal{P}_{d}(\mathfrak{gl}_{d}(A))$ and $S_{d}\wr A$-mod cannot be equivalent in this case.
@@
This is an equivalence of categories whenever $n\geq d$. See [12] for details.
By [11, Lemma 11.1.1], there is a surjective algebra map

### v2

@@ 3.3. Defining a $S_{d}\wr A$-module structure
be the function which counts the number of entries of $(\mathbf{i},\mathbf{i})_{k}$ that are different than the corresponding entries of $((1,2,\dotsc,d),(1,2,\dotsc,d))_{k}$. We argue by induction on $D\left((\mathbf{i},\mathbf{i})_{k}\right)$, with the base case of $D\left((\mathbf{i},\mathbf{i})_{k}\right)=0$ being trivial.
Assume $D\left((\mathbf{i},\mathbf{i})_{k}\right)>0$. Then $i_{\ell}\neq\ell$ for some $\ell\in[1,d]$. We first consider the case when $\ell$ does not appear as an entry of $(\mathbf{i},\mathbf{i})_{k}$. If $\ell\neq k$, then we may use (3) from the previous lemma to simultaneously change $i_{\ell}$ of both tuples to $\ell$ and thereby obtain a pair $(\mathbf{i}^{\prime},\mathbf{i}^{\prime})_{k}\in\tilde{I}^{2}(n,d)$ which is equivalent to $(\mathbf{i},\mathbf{i})_{k}$ and with $D((\mathbf{i}^{\prime},\mathbf{i}^{\prime})_{k})=D((\mathbf{i},\mathbf{i})_{k})-1$. If $\ell=k$, let $\mathbf{i}^{\prime}=(i_{1},\dotsc,i_{k-1},\ell,i_{k+1},\dotsc,i_{d})$. Using (1) shows that $\alpha^{a}_{(\mathbf{i},\mathbf{i})_{k}}=\alpha^{a}_{(\mathbf{i}^{\prime},\mathbf{i})_{k}}$ and (2) shows $\alpha^{a}_{(\mathbf{i}^{\prime},\mathbf{i})_{k}}=\alpha^{a}_{(\mathbf{i}^{\prime},\mathbf{i}^{\prime})_{k}}$. Once again, we have obtained a pair $(\mathbf{i}^{\prime},\mathbf{i}^{\prime})_{k}\in\tilde{I}^{2}(n,d)$ which is equivalent to $(\mathbf{i},\mathbf{i})_{k}$ and with $D((\mathbf{i}^{\prime},\mathbf{i}^{\prime})_{k})=D((\mathbf{i},\mathbf{i})_{k})-1$.
@@
By induction, $(\mathbf{i},\mathbf{i})_{k}$ is equivalent to $((1,2,\dotsc,d),(1,2,\dotsc,d))_{k}$ and this proves the claim. ∎
In light of the previous result, we sometimes write $\alpha^{a}_{k}$ for $\alpha^{a}_{(\mathbf{j},\mathbf{i})_{k}}$, where the $(\mathbf{j},\mathbf{i})_{k}\in\tilde{I}^{2}(n,d)$ is left implicit.
We continue to assume that $N$ is a $\mathds{k}S_{d}$-supermodule such that $E_{n}^{\otimes d}\otimes_{\mathds{k}S_{d}}N$ has a $U(\mathfrak{gl}_{n}(A))$-supermodule structure that makes it a supermodule in $\mathcal{P}_{d}(\mathfrak{gl}_{n}(A))$, and such that the action of the subalgebra $U(\mathfrak{gl}_{n}(\mathds{k}))$ on $E_{n}^{\otimes d}\otimes_{\mathds{k}S_{d}}N$ is the obvious one.
@@
We first observe that $\alpha_{k}^{1}:=\alpha_{k}^{1_{A}}$ is the identity map for all $k\in[1,d]$. Namely, fix $(\mathbf{j},\mathbf{i})_{k}\in\tilde{I}^{2}(n,d)$ and then
@@
From Proposition 3.1.1, $\alpha_{k}^{1}(n)=n$, as claimed.
@@
for all $k\in[1,d]$ and all homogeneous $a\in A$. Since the various $1^{\otimes(k-1)}\otimes a\otimes 1^{\otimes(d-k)}$ generate $A^{\otimes d}$, it suffices to verify:
@@
We next explain the fourth claim. Pick $(\mathbf{j},\mathbf{i})_{k},(\mathbf{i},\mathbf{i})_{k}\in\tilde{I}^{2}(n,d)$ with $i_{k}$ and $j_{k}$ distinct. Two such pairs exist because $n\geq d+1$. We wish to show $\alpha_{(\mathbf{j},\mathbf{i})_{k}}^{a}\circ\alpha_{(\mathbf{i},\mathbf{i})_{k}}^{a^{\prime}}=\alpha_{(\mathbf{j},\mathbf{i})_{k}}^{aa^{\prime}}$. Using the supercommutator formula (2.2.1), we have
@@
However, by the second property of supermodules in $\mathcal{P}_{d}(\mathfrak{gl}_{n}(A))$ the first term in the last sum equals zero. Proposition 3.1.1 shows that $\alpha_{(\mathbf{j},\mathbf{i})_{k}}^{a}(\alpha_{(\mathbf{i},\mathbf{i})_{k}}^{a^{\prime}}(n))=\alpha_{(\mathbf{j},\mathbf{i})_{k}}^{aa^{\prime}}(n)$, as desired.
@@
Fix $(\mathbf{j},\mathbf{i})_{p}\in\tilde{I}^{2}(n,d)$. Set $\mathbf{j}^{\prime}=(j_{1},\dotsc j_{p-1},j_{p+1},j_{p},j_{p+2},\dotsc,j_{d})$ and $\mathbf{i}^{\prime}=(i_{1},\dotsc i_{p-1},i_{p+1},i_{p},i_{p+2},\dotsc,i_{d})$. Then $(\mathbf{j}^{\prime},\mathbf{i}^{\prime})_{p+1}\in\tilde{I}^{2}(n,d)$. Under the place permutation action of $S_{d}$ on $E_{n}^{\otimes d}$, we have $\mathbf{v}_{\mathbf{j}^{\prime}}\sigma_{p}=\mathbf{v}_{\mathbf{j}}$ and $\mathbf{v}_{\mathbf{i}}\sigma_{p}=\mathbf{v}_{\mathbf{i}^{\prime}}$. We then calculate:
@@
Proposition 3.1.1 implies $\sigma_{p}(\alpha_{p}^{a}(\sigma_{p}(n)))=\alpha^{a}_{p+1}(n)$, as desired. ∎
@@ 3.4. Recovering the action of $\mathfrak{gl}_{n}(A)$
We continue to assume that $N$ is a $\mathds{k}S_{d}$-supermodule such that $E_{n}^{\otimes d}\otimes_{\mathds{k}S_{d}}N$ has a $U(\mathfrak{gl}_{n}(A))$-supermodule structure that makes it a supermodule in $\mathcal{P}_{d}(\mathfrak{gl}_{n}(A))$, and such that the action of the subalgebra $U(\mathfrak{gl}_{n}(\mathds{k}))$ on $E_{n}^{\otimes d}\otimes_{\mathds{k}S_{d}}N$ is the obvious one. In the previous section, the action of $U(\mathfrak{gl}_{n}(A))$ on $E_{n}^{\otimes d}\otimes_{\mathds{k}S_{d}}N$ was used on certain special elements to define maps $\alpha^{a}_{k}:N\to N$ for homogeneous $a\in A$ and $k\in[1,d]$. We now show that these maps completely determine the action of $U(\mathfrak{gl}_{n}(A))$ on $E_{n}^{\otimes d}\otimes_{\mathds{k}S_{d}}N$.
@@
Let $n\geq d+1$ and let $E_{n}^{\otimes d}\otimes_{\mathds{k}S_{d}}N$ be as above and let $n\in N$. For homogeneous $a\in A$, $p,q\in[1,n]$, and $\mathbf{i}\in I(n,d)$, we have
@@
The base case is when all entries of $\mathbf{i}$ are distinct and $D(\mathbf{i})=0$. If $q$ does not appear as an entry of $\mathbf{i}$, then the left side of (3.4.1) is zero because $\boldsymbol{\varepsilon}_{\mathbf{i}}+\varepsilon_{p}-\varepsilon_{q}$ is not an element of $\Lambda(n,d)$, and the right side is also zero because $\delta_{q,i_{k}}=0$ for all $k$. If $q$ appears as an entry of $\mathbf{i}$, then it appears only once — say at position $\ell\in[1,d]$. Let $\mathbf{j}=(i_{1},\dotsc,i_{\ell-1},p,i_{\ell+1},\dotsc,d)$. In this case, the left hand side equals $\mathbf{v}_{\mathbf{j}}\otimes\alpha^{a}_{(\mathbf{j},\mathbf{i})_{\ell}}(n)=\mathbf{v}_{\mathbf{j}}\otimes\alpha^{a}_{\ell}(n)$ and, by direct calculation the right hand side also equals $\mathbf{v}_{\mathbf{j}}\otimes\alpha^{a}_{\ell}(n)$. In both cases (3.4.1) is verified when $D(\mathbf{i})=0$ for all $E^{a}_{p,q}$.
We now assume $D(\mathbf{i})>0$ and that (3.4.1) holds for all $E^{a}_{p,q}$ and all $v_{\mathbf{j}}\otimes n$ whenever $D(\mathbf{j})<D(\mathbf{i})$. Since $D(\mathbf{i})>0$, there exists $1\leq k_{1}<k_{2}\leq d$ such that $i_{k_{1}}=i_{k_{2}}$. Let $\ell=k_{1}$. Since $n\geq d+1$, we may choose $x\in[1,n]$ so that $x$ does not appear as an entry of $\mathbf{i}$. Set $\mathbf{j}=(i_{1},\dotsc,i_{\ell-1},x,i_{\ell+1},\dotsc,i_{d})$. Observe that $D(\mathbf{j})<D(\mathbf{i})$. From the supercommutator formula (2.2.1), we have
@@ 3.6. Main Theorem
If $n\geq d+1$ and $N$ is a $S_{d}\wr A$-supermodule, then $F^{A}_{n,d}(N)\cong E_{n}^{\otimes d}\otimes_{\mathds{k}S_{d}}N$ is finite-dimensional if and only if $N$ is finite-dimensional. Thus, the restriction of $F^{A}_{n,d}$ defines an equivalence between the full subcategories of finite-dimensional supermodules.
@@ 3.7. Some Consequences
We give a few example applications of the main theorem. For simplicity’s sake, throughout this section we assume $A$ is finite-dimensional and that $\mathds{k}$ is a splitting field for $A$ (e.g., $\mathds{k}$ is algebraically closed).
@@
Write $J(A)$ for the Jacobson radical of the superalgebra $A$. By [Brundan and Kleshchev, 2002, Lemma 2.6] this is a homogeneous ideal and coincides with the Jacobson radical of $A$ as an algebra. The arguments of [Chuang and Tan, 2003, Section 3] show that
@@
By [Chuang and Tan, 2003, Section 6], if $A$ is quasi-hereditary algebra, then so is $S_{d}\wr A$ and, hence, the category of $S_{d}\wr A$-modules is a highest weight category. Thus, we have the following result.
@@
Say $A$ is a finite-dimensional $\mathds{k}$-superalgebra and that $S=\left\{S_{1},\dotsc,S_{t}\right\}$ is a complete, irredundant set of simple $A$-supermodules up to isomorphism. Recall that a simple $A$-supermodule may either be of type M or of type Q. It is known that the simple $S_{d}\wr A$-supermodules are indexed by the set
@@
For example, see Macdonald [1980] for the construction of the simples when $A$ is an algebra, and see Rosso and Savage [2017] or Davidson et al. [2023] for the construction of the simples when $A$ is a superalgebra. Write $S^{\pi}$ for the simple $S_{d}\wr A$-supermodule labelled by $\pi\in\Pi^{d}(A)$. We note that the type of $S^{\pi}$ can be determined from the function $\pi$ and that $F^{A}_{n,d}$ preserves the type of simple supermodules. Since we do not use this information, we omit the details.
@@
$\left\{L_{n}^{A}(\pi):=F^{A}_{n,d}(S^{\pi})\mid\pi\in\Pi^{d}(A)\right\}$
@@ 3.8. Sharpness of the Bound
Given that the equivalence of categories holds when $n\geq d$ in the classical (i.e., $A=\mathds{k}$) case, we could ask if the bound $n\geq d+1$ here is an artifact of the proof. The following example shows this is not the case. A similar observation about the bound was made for $A=\mathds{k}[t^{\pm 1}]$ in Flicker [2021].
@@
Let $d=n$, and let $A=\mathds{k}[x]/(x^{2}-1)$ be the group algebra of the cyclic group of order two. We use the notation established in Example 2.4.3.
Let $L=\mathds{k}u$ be a one-dimensional vector space viewed as a $U(\mathfrak{gl}_{n}(A))$-module by declaring that $x.u=0$ for $x\in\mathfrak{sl}_{n}(A)$, $I_{n}.u=nu$, and $xI_{n}.u=u$. Observe that $L$ is a polynomial representation of $U(\mathfrak{gl}_{n}(\mathds{k}))$ of degree $d=n$. The spanning vector has weight $\boldsymbol{\varepsilon}_{\mathbf{i}}=\varepsilon_{1}+\dotsb+\varepsilon_{n}$ and, since $\boldsymbol{\varepsilon}_{\mathbf{i}}(h_{i})=1$ for $i=1,\dotsc,n$, this module vacuously satisfies the second condition of Definition 2.4.1. Thus, $L$ is in $\mathcal{P}_{d}(\mathfrak{gl}_{n}(A))$. In fact, it is not difficult to see that $L$ is in the image of $F^{A}_{n,d}$.
@@
Observe that $T$ restricted to $U(\mathfrak{gl}_{n}(\mathds{k}))$ is a polynomial representation of degree $d=n$. It again vacuously satisfies the second condition of Definition 2.4.1. Therefore, $T$ is in $\mathcal{P}_{d}(\mathfrak{gl}_{d}(A))$ and is a non-split extension of $L$ with itself. On the other hand, $S_{d}\wr A$ is a semisimple algebra. Therefore, $\mathcal{P}_{d}(\mathfrak{gl}_{d}(A))$ and $S_{d}\wr A$-mod cannot be equivalent in this case.
@@
This is an equivalence of categories whenever $n\geq d$. See [Evseev and Kleshchev, 2017, Section 5] for details.
By [Davidson et al., 2023, Lemma 11.1.1], there is a surjective algebra map